OpenStudy (anonymous):

Please help. is there a way to find the four solutions for 1/2 + 1/2cos(2x) =a given that 0

OpenStudy (anonymous):

If you pick a value for x lets say pi/4 Then you can find a value for a which would be 1/2+1/2cos(2*pi/4) = a So then a = 1/2 and that works, So now you just have to do that for more values of x and those give you your values of a. You just need to make sure a is between 0 and 1

OpenStudy (jdoe0001):

is there more to it? more info?

OpenStudy (jdoe0001):

looks pretty ambiguous to me

OpenStudy (anonymous):

@Arfney thanks, what can I do after cause when we move the1/2 to it the equation become 0

OpenStudy (anonymous):

@jdoe0001 nope,that is all I got T T

OpenStudy (jdoe0001):

hmm

OpenStudy (anonymous):

In my method you "guess" x and solve for A. Then if A is between 0 and 1 then you have a solution.

OpenStudy (anonymous):

@Arfney thanks again :)

OpenStudy (anonymous):

Your welcome! :D

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{2}+\cfrac{1}{2}cos(2x)=a\implies \cfrac{1}{2}+\cfrac{cos(2x)}{2}=a\implies \cfrac{1+cos(2x)}{2}=a \\ \quad \\ 1+cos(2x)=2a\implies \cancel{ 1 }+[{\color{brown}{ 2cos^2(x)\cancel{ -1}}}]=2a \\ \quad \\ 2cos^2(x)=2a\implies cos^2(x)=a\implies cos(x)=\pm \sqrt{a} \\ \quad \\ x=cos^{-1}(\pm \sqrt{a})\) is thus far what I have

OpenStudy (jdoe0001):

but I can see many "a" between 0 and 1, having an angle

OpenStudy (anonymous):

@jdoe0001 yeah I can see many a too :( and thanks so much

OpenStudy (jdoe0001):

hmmm unless you're expected some rational number..... I gather... as Arfney said, if you use say 1/4 for that then \(\bf x=cos^{-1}\left(\pm \sqrt{\frac{1}{4}}\right)\implies x=cos^{-1}\left(\pm \sqrt{\frac{1^2}{2^2}}\right)\implies x=cos^{-1}\left(\pm \frac{1}{2}\right)\) and that should give you an angle, and then I guess you could use the reference angle of that for the other 3 quadrants so that angle in each quadrant, would give 4 angles that are between \(\bf 0..2\pi\)